Jaydens brother is 7 years older than jayden. His mother is 4 less than 3 times jaydens age. Their combined age is 103.How old is jayden his brother and his mother
Answer by VFBundy(438)

Jayden's age = j
Jayden's brother's age = j + 7
Jayden's mother's age = 3j - 4

j + (j + 7) + (3j - 4) = 103

j + j + 7 + 3j - 4 = 103

5j + 3 = 103

5j = 100

j = 20

Jayden's age = j = 20
Jayden's brother's age = j + 7 = 20 + 7 = 27
Jayden's mother's age = 3j - 4 = 3(20) - 4 = 60 - 4 = 56